One number is fourteen more than twice another. If their sum is decreased by nine, the result is seventeen. Find the numbers. (Enter solutions from smallest to largest.)
Answer by checkley77(12844)

x=2y+14
x+y-9=14
x=y=14+9
x+y=23
x=-y+23 multiply by 2 & add these 2 equations.
x=2y+14
2x=-2y+46
------------------
3x=60
x=60/3
x=20 ans.
20=2y+14
20-14=2y
2y=6
y=6/2
y=3 ans.